OpenStudy (anonymous):

448÷〖(-2)〗^3÷2×(-3)+1/5 HELPPPPP

OpenStudy (anonymous):

Hi, I will help you

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

-2^ 3 = 8 / -6 = 4/ -3 448 / 4/ -3 = 5376 + 1/5 = 26881/ 5

OpenStudy (anonymous):

i forgot to put -2^3 = -8 not 8

OpenStudy (anonymous):

but didnt change answer

OpenStudy (anonymous):

so i divde 26681 by 5 to get the answer

OpenStudy (anonymous):

yes or leave it like that

OpenStudy (anonymous):

5376

OpenStudy (anonymous):

thats not right

OpenStudy (anonymous):

i got 7